The drama club sold 800 tickets to their spring play. Floor seats cost $10 each & balcony seats cost $8 each. Ticket sales were $7440. How many floor seat tickets were sold?






please help me!

Answer by gems(12)

Let x be the number of tickets for Floor seats
y be the number of tickets for Balcony seats
Equation:
x + y = 800
10x + 8y = 7440
By substitution method
x + y = 800 or x = 800 - y
Substitute to 10x + 8y = 7440
10(800-y) + 8y = 7440
8000 - 10y + 8y = 7440
-2y = 7400 - 8000
-2y = - 560
y = 280
If y = 280 then x = 800 - y, or x = 800 - 280
then x = 520
So there were 520 Floor seats sold